# On second-order variational analysis of variational convexity of prox-regular functions

### v1

@@ 4 On neighborhood-based second-order characterizations of variational $s$-convexity
Next we prove that (iv) is equivalent with (v). Obviously, (v) implies (iv). Now assume that (iv) holds with some open convex neighborhood $U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ and some $\rho>f(\bar{x})$ and consider at a point $(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)$ and a subspace $L\in{\mathcal{S}}_{f}\partial f(x,x^{*})$ together with a sequence $(x_{k},x_{k}^{*})\subset{\cal O}^{f}_{\partial f}$ satisfying $(x_{k},x_{k}^{*})\mathop{\longrightarrow}\limits_{f}^{\mathrm{gph}\,\partial f}(x,x^{*})$ and $d_{{\cal Z}}(L,L_{k})\to 0$, where $L_{k}:=T^{f}_{\mathrm{gph}\,\partial f}(x_{k},x_{k}^{*})$. Denoting by $P$ and $P_{k}$ the orthogonal projections onto $L$ and $L_{k}$, respectively, we obtain that for every $(z,z^{*})\in L$ the sequence $(z_{k},z_{k}^{*}):=P_{k}(z,z^{*})\in L_{k}$ converges to $P(z,z^{*})=(z,z^{*})$. For all $k$ sufficiently large we have $(x_{k},x_{k}^{*})\in{\cal O}^{f}_{\partial f}\cap((U\cap L_{f}(\rho))\times V)$ and therefore $\langle z_{k}^{*},z_{k}\rangle\geq 0$. This implies $\langle z^{*},z\rangle=\lim_{k\to\infty}\langle z_{k}^{*},z_{k}\rangle\geq 0$ and the implication (iv)$\Rightarrow$(v) is established.
@@
In order to show the equivalence between (v) and (vi), note that for $(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)$ and $L={\rm rge\;}(P,W)\in{\mathcal{S}}_{f}(\partial f(x,x^{*})$ with $(P,W)\in{\cal M}_{P,W}^{f}\partial f(x,x^{*})$ the condition $\langle z^{*},z\rangle\geq 0$, $(z,z^{*})\in L$ is equivalent with
@@ 5 Point-based second-order characterizations of variational strong convexity and tilt-stability
ad ”(i)$\Rightarrow$(ii)”: The statement follows from [16, Theorem 3] together with the fact that variational convexity implies prox-regularity, cf. Remark 2.8(i).
@@
Consider $x^{*}\in{\cal B}_{\epsilon^{\prime}}(0)$. Since $(x^{*},M_{\gamma}(x^{*}))\in\mathrm{gph}\,\partial g$ we have $(M_{\gamma}(x^{*}),x^{*})\in\mathrm{gph}\,\partial\psi$ by [15, Theorem 23.5] and, since $\psi$ is an lsc convex function, $M_{\gamma}(x^{*})\in\mathop{\rm arg\,min}\{\psi-\langle x^{*},\cdot\rangle\}$ follows. Actually $M_{\gamma}(x^{*})$ is the unique minimizer of $\psi-\langle x^{*},\cdot\rangle$, since any minimizer $\tilde{x}$ of $\psi-\langle x^{*},\cdot\rangle$ satisfies $x^{*}\in\partial\psi(\tilde{x})$ and therefore $\tilde{x}\in\partial g(x^{*})=\{M_{\gamma}(x^{*})\}$ by [15, Theorem 23.5]. Thus we conclude that $\bar{x}$ is a tilt-stable minimizer for $\psi$ with modulus $\kappa$ and ${\rm tilt\,}(\psi;\bar{x})={\rm tilt\,}(f;\bar{x})$.
Consider a point $(x,x^{*})\in\mathrm{gph}\,\partial\psi\cap({\cal B}_{\gamma}(\bar{x})\times{\cal B}_{\epsilon^{\prime}}(0))$. Then $(x^{*},x)\in\mathrm{gph}\,\partial g\cap({\cal B}_{\epsilon^{\prime}}(0)\times{\cal B}_{\gamma}(\bar{x}))$ and $g(x^{*})+\psi(x)=\langle x^{*},x\rangle$ by [15, Theorem 23.5] implying $x=\nabla g(x^{*})=M_{\gamma}(x^{*})$ and $\psi(x)=\langle x^{*},x\rangle-(\langle x^{*},M_{\gamma}(x^{*})\rangle-f(M_{\gamma}(x^{*})))=f(x)$. Further, by (5.21) together with $x=M_{\gamma}(x^{*})\in{\cal B}_{\gamma}(\bar{x})$ the first-order optimality condition
@@
Since $\psi$ is a lsc convex function, $\partial\psi$ is a maximally monotone mapping and so is $rI+\partial\psi$ as well. Therefore, the mapping $S:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ given by $\mathrm{gph}\,S=\mathrm{gph}\,(rI+\partial\psi)\cap(\tilde{U}\times\tilde{V})$ is locally maximal monotone around $(\bar{x},\rho\bar{x})$ by Proposition 2.3. By the proof of [20, Theorem 13.36], the mapping $rI+\mathcal{T}^{f}_{\epsilon}$ is monotone and, since $\mathrm{gph}\,S\subset\mathrm{gph}\,(rI+\mathcal{T}^{f}_{\epsilon})$, there exists neighborhoods $\hat{U}$ of $\bar{x}$ and $\hat{V}$ of $r\bar{x}$ such that
@@
Hence $D^{*}(rI+\partial\psi)(\bar{x},r\bar{x})=D^{*}(rI+\mathcal{T}^{f}_{\epsilon})(\bar{x},r\bar{x})$ and by elementary calculus rules for coderivatives, cf. [20, Exercise 10.43], we obtain that $D^{*}(\partial\psi)(\bar{x},0)=D^{*}\mathcal{T}^{f}_{\epsilon}(\bar{x},0)=D^{*}_{f}(\partial f)(\bar{x},0)$, where the last equality follows from Proposition 3.5. The lsc convex function $\psi$ is prox-regular and subdifferential continuous at $\bar{x}$ for 00 and we can use Theorem 2.12 to obtain
@@
ad ”(v)$\Rightarrow$(i)”: Let $r\geq 0$ and $\epsilon>0$ denote the parameter values for prox-regularity of $f$ at $\bar{x}$ for 00. By Remark 2.8, $f$ is variationally $(-r)$-convex and let $\widehat{f},U,V$ and $\rho$ be as Definition 2.6. By possibly lowering $\epsilon$ we may assume that $f(\bar{x})+\epsilon\leq\rho$ and then, by possibly shrinking $U\times V$ and taking into account subdifferential continuity of $\widehat{f}$ at $\bar{x}$ for ${\bar{x}^{\ast}}$, we may assume that $U\times V\subset{\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}(0)$ and $\widehat{f}(x)<f(\bar{x})+\epsilon$, $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ implying $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)=\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$. Hence ${\mathcal{S}}(\partial\widehat{f})(\bar{x},0)={\mathcal{S}}_{f}(\partial f)(\bar{x},0)$ by Proposition 3.5 implying ${\cal M}_{P,W}\partial\widehat{f}(\bar{x},0)={\cal M}_{P,W}^{f}\partial f(\bar{x},0)$.
Assume now that (v) holds, pick $\sigma^{\prime}\in(0,\sigma)$ and consider the function $\tilde{f}:=\widehat{f}-\frac{\sigma^{\prime}}{2}\|\cdot-\bar{x}\|^{2}$. Since $\widehat{f}$ is prox-regular and subdifferentially continuous at $\bar{x}$ for 00, $\tilde{f}$ has these properties as well. By [3, Proposition 3.15] we have
@@
From the proof of this theorem we can also extract the following corollary.
Corollary 5.2.
@@
The implication (i)$\Rightarrow$(ii) follows from [16, Theorem 3] together with the fact that variational convexity implies prox-regularity, cf. Remark 2.8(i). In order to show the reverse implication, consider the lsc convex function $\psi$ from the proof of the implication (ii)$\Rightarrow$(iii) in Theorem 5.1 with $\sigma^{\prime}$ replaced by $\sigma$. Then $\psi$ is prox-regular and subdifferentially continuous, and since $\bar{x}$ is a tilt-stable minimizer for $\psi$ with modulus $\sigma^{-1}$, $\psi$ is variationally strongly convex at $\bar{x}$ for ${\bar{x}^{\ast}}$ with modulus $\sigma$ by Theorem 2.13 and the same property follows for $f$ by Lemma 4.1. ∎
@@
    For every $s^{\prime}\in(0,s)$, $f$ is variationally $s^{\prime}$-convex at $\bar{x}$ for ${\bar{x}^{\ast}}$.
@@
Choose $t>0$ such that $\sigma:=s+t>0$ and apply Theorem 5.1 to the function $\tilde{f}:=f+\frac{t}{2}\|\cdot-\bar{x}\|^{2}$. Then the assertions follow from Proposition 3.2 and Lemma 4.2. ∎

### v2

@@ 4 On neighborhood-based second-order characterizations of variational $s$-convexity
Next we prove that (iv) is equivalent with (v). Obviously, (v) implies (iv). Now assume that (iv) holds with some open convex neighborhood $U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ and some $\rho>f(\bar{x})$ and consider at a point $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)$ and a subspace $L\in{\mathcal{S}}_{f}\partial f(x,x^{*})$ together with a sequence $(x_{k},x_{k}^{*})\subset{\cal O}^{f}_{\partial f}$ satisfying $(x_{k},x_{k}^{*})\mathop{\longrightarrow}\limits_{f}^{\mathrm{gph}\,\partial f}(x,x^{*})$ and $d_{{\cal Z}}(L,L_{k})\to 0$, where $L_{k}:=T^{f}_{\mathrm{gph}\,\partial f}(x_{k},x_{k}^{*})$. Denoting by $P$ and $P_{k}$ the orthogonal projections onto $L$ and $L_{k}$, respectively, we obtain that for every $(z,z^{*})\in L$ the sequence $(z_{k},z_{k}^{*}):=P_{k}(z,z^{*})\in L_{k}$ converges to $P(z,z^{*})=(z,z^{*})$. For all $k$ sufficiently large we have $(x_{k},x_{k}^{*})\in{\cal O}^{f}_{\partial f}\cap((U\cap L_{f}(\rho))\times V)$ and therefore $\langle z_{k}^{*},z_{k}\rangle\geq 0$. This implies $\langle z^{*},z\rangle=\lim_{k\to\infty}\langle z_{k}^{*},z_{k}\rangle\geq 0$ and the implication (iv)$\Rightarrow$(v) is established.
@@
In order to show the equivalence between (v) and (vi), note that for $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)$ and $L={\rm rge\;}(P,W)\in{\mathcal{S}}_{f}(\partial f(x,x^{*})$ with $(P,W)\in{\cal M}_{P,W}^{f}\partial f(x,x^{*})$ the condition $\langle z^{*},z\rangle\geq 0$, $(z,z^{*})\in L$ is equivalent with
@@
Remark 4.4.
Very recently, Khanh, Khoa, Mordukhovich and Phat [11, Theorem 5.5] gave a neighborhood-based characterization of variational convexity in terms of regular coderivatives as well as limiting coderivatives of the $f$-attentive $\epsilon$-localization of $\partial f$ around $(\bar{x},{\bar{x}^{\ast}})$. By taking into account Proposition 3.5 and Lemma 4.2, the equivalences (i)$\Leftrightarrow$(ii)$\Leftrightarrow$(iii) in Theorem 4.3 can also be derived from this result.
@@ 5 Point-based second-order characterizations of variational strong convexity and tilt-stability
ad ”(i)$\Rightarrow$(ii)”: The statement follows from [18, Theorem 3] together with the fact that variational convexity implies prox-regularity, cf. Remark 2.8(i).
@@
Consider $x^{*}\in{\cal B}_{\epsilon^{\prime}}(0)$. Since $(x^{*},M_{\gamma}(x^{*}))\in\mathrm{gph}\,\partial g$ we have $(M_{\gamma}(x^{*}),x^{*})\in\mathrm{gph}\,\partial\psi$ by [17, Theorem 23.5] and, since $\psi$ is an lsc convex function, $M_{\gamma}(x^{*})\in\mathop{\rm arg\,min}\{\psi-\langle x^{*},\cdot\rangle\}$ follows. Actually $M_{\gamma}(x^{*})$ is the unique minimizer of $\psi-\langle x^{*},\cdot\rangle$, since any minimizer $\tilde{x}$ of $\psi-\langle x^{*},\cdot\rangle$ satisfies $x^{*}\in\partial\psi(\tilde{x})$ and therefore $\tilde{x}\in\partial g(x^{*})=\{M_{\gamma}(x^{*})\}$ by [17, Theorem 23.5]. Thus we conclude that $\bar{x}$ is a tilt-stable minimizer for $\psi$ with modulus $\kappa$ and ${\rm tilt\,}(\psi;\bar{x})={\rm tilt\,}(f;\bar{x})$.
Consider a point $(x,x^{*})\in\mathrm{gph}\,\partial\psi\cap({\cal B}_{\gamma}(\bar{x})\times{\cal B}_{\epsilon^{\prime}}(0))$. Then $(x^{*},x)\in\mathrm{gph}\,\partial g\cap({\cal B}_{\epsilon^{\prime}}(0)\times{\cal B}_{\gamma}(\bar{x}))$ and $g(x^{*})+\psi(x)=\langle x^{*},x\rangle$ by [17, Theorem 23.5] implying $x=\nabla g(x^{*})=M_{\gamma}(x^{*})$ and $\psi(x)=\langle x^{*},x\rangle-(\langle x^{*},M_{\gamma}(x^{*})\rangle-f(M_{\gamma}(x^{*})))=f(x)$. Further, by (5.21) together with $x=M_{\gamma}(x^{*})\in{\cal B}_{\gamma}(\bar{x})$ the first-order optimality condition
@@
Since $\psi$ is a lsc convex function, $\partial\psi$ is a maximally monotone mapping and so is $rI+\partial\psi$ as well. Therefore, the mapping $S:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ given by $\mathrm{gph}\,S=\mathrm{gph}\,(rI+\partial\psi)\cap(\tilde{U}\times\tilde{V})$ is locally maximal monotone around $(\bar{x},\rho\bar{x})$ by Proposition 2.3. By the proof of [22, Theorem 13.36], the mapping $rI+\mathcal{T}^{f}_{\epsilon}$ is monotone and, since $\mathrm{gph}\,S\subset\mathrm{gph}\,(rI+\mathcal{T}^{f}_{\epsilon})$, there exist neighborhoods $\hat{U}$ of $\bar{x}$ and $\hat{V}$ of $r\bar{x}$ such that
@@
Next we choose some neighborhoods $\bar{U}$ of $\bar{x}$ and $\bar{V}$ of 00 satisfying $\bar{U}\subset\tilde{U}\cap\hat{U}$ and $r\bar{U}+\bar{V}\subset\tilde{V}\cap\hat{V}$. It follows that $\mathrm{gph}\,\partial\psi\cap(\bar{U}\times\bar{V})=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(\bar{U}\cap\bar{V})$ and thus we obtain $D^{*}(\partial\psi)(\bar{x},0)=D^{*}\mathcal{T}^{f}_{\epsilon}(\bar{x},0)=D^{*}_{f}(\partial f)(\bar{x},0)$, where the last equality follows from Proposition 3.5. The lsc convex function $\psi$ is prox-regular and subdifferential continuous at $\bar{x}$ for 00 and we can use Theorem 2.12 to obtain
@@
ad ”(v)$\Rightarrow$(i)”: Let $r\geq 0$ and $\epsilon>0$ denote the parameter values for prox-regularity of $f$ at $\bar{x}$ for 00. By Remark 2.8, $f$ is variationally $(-r)$-convex and let $\widehat{f},U,V$ and $\rho$ be as Definition 2.6. By possibly lowering $\epsilon$ we may assume that $f(\bar{x})+\epsilon\leq\rho$ and then, by possibly shrinking $U\times V$ and taking into account subdifferential continuity of $\widehat{f}$ at $\bar{x}$ for ${\bar{x}^{\ast}}$, we may assume that $U\times V\subset{\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}(0)$ and $\widehat{f}(x)<f(\bar{x})+\epsilon$, $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)$ implying $\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)=\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)=\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}\cap(U\times V)$. Hence ${\mathcal{S}}(\partial\widehat{f})(\bar{x},0)={\mathcal{S}}_{f}(\partial f)(\bar{x},0)$ by Proposition 3.5 implying ${\cal M}_{P,W}\partial\widehat{f}(\bar{x},0)={\cal M}_{P,W}^{f}\partial f(\bar{x},0)$.
Assume now that (v) holds, pick $\sigma^{\prime}\in(0,\sigma)$ and consider the function $\tilde{f}:=\widehat{f}-\frac{\sigma^{\prime}}{2}\|\cdot-\bar{x}\|^{2}$. Since $\widehat{f}$ is prox-regular and subdifferentially continuous at $\bar{x}$ for 00, $\tilde{f}$ has these properties as well. By [4, Proposition 3.15] we have
@@
We are now in the position to extend the equivalences of Theorem 2.7 for variationally strongly convex functions.
Theorem 5.2.
@@
  3. (iii)
    $\bar{x}$ is a local minimizer for $f-\langle{\bar{x}^{\ast}},\cdot\rangle$ and there is some $\rho>f(\bar{x})$ such that the $f$-truncation $\partial_{\rho}f$ is strongly metrically regular at $\bar{x}$ for ${\bar{x}^{\ast}}$ with modulus $\sigma^{-1}$.
@@
In order to unburden notation, we may again assume that ${\bar{x}^{\ast}}=0$. The implication (i)$\Rightarrow$(ii) follows from [18, Theorem 3] together with the fact that variational convexity implies prox-regularity, cf. Remark 2.8(i). In order to show the reverse implication, consider the lsc convex function $\psi$ from the proof of the implication (ii)$\Rightarrow$(iii) in Theorem 5.1 with $\sigma^{\prime}$ replaced by $\sigma$. Then $\psi$ is prox-regular and subdifferentially continuous, and since $\bar{x}$ is a tilt-stable minimizer for $\psi$ with modulus $\sigma^{-1}$, $\psi$ is variationally strongly convex at $\bar{x}$ for ${\bar{x}^{\ast}}$ with modulus $\sigma$ by Theorem 2.13 and the same property follows for $f$ by Lemma 4.1.
For showing the implication (ii)$\Rightarrow$(iii), we again resort to the proof of the implication (ii)$\Rightarrow$(iii) in Theorem 5.1. From tilt-stability of $\bar{x}$ for $\psi$ with modulus $\sigma^{-1}$ it follows from Theorem 2.12 that $\psi$ is metrically strongly regular at $\bar{x}$ with the same modulus. Since we have also shown that $\mathrm{gph}\,\partial\psi$ and $\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}$ coincide locally around $(\bar{x},0)$, $\mathcal{T}^{f}_{\epsilon}$ is also strongly metrically regular at $\bar{x}$ for 00 with modulus $\sigma^{-1}$. By definition, $\mathcal{T}^{f}_{\epsilon}$ is a localization of the $f$-attentive truncation $\partial_{f(\bar{x})+\epsilon}f$ and thus statement (iii) follows with $\rho=f(\bar{x})+\epsilon$.
Finally let us show the reverse direction (iii)$\Rightarrow$(ii). Assume that (iii) holds, let $f$ be prox-regular with parameters $r\geq 0$ and $\epsilon>0$ and consider a real $\rho>f(\bar{x})$, open neighborhoods $U$ of $\bar{x}$ and $V$ of 00 together with a mapping $\vartheta:V\to U$ which is Lipschitz continuous with constant $\sigma^{-1}$, such that $\mathrm{gph}\,(\partial_{\rho}f)^{-1}\cap(V\times U)=\mathrm{gph}\,\vartheta$. We claim that $\bar{x}$ is a strict local minimizer of $f$, i.e., there exists some $\gamma>0$ such that $f(x)>f(\bar{x})$ holds for all $x\in\bar{\cal B}_{\gamma}(\bar{x})\setminus\{\bar{x}\}$. Assuming on the contrary that this claim does not hold, we can find a sequence $x_{k}\to\bar{x}$, $x_{k}\not=\bar{x}$, satisfying $f(x_{k})\leq f(\bar{x})$. Since $\bar{x}$ is a local minimizer of $f$, we can conclude that $f(x_{k})=f(\bar{x})$ for all $k$ sufficiently large and consequently $x_{k}$ is a local minimizer for $f$ as well. Hence, $0\in\partial f(x_{k})=\partial_{\rho}f(x_{k})$ contradicting our assumption that $(\partial_{\rho}f)^{-1}$ has a single-valued localization around $(0,\bar{x})$. Hence, our claim holds true and, since $f$ is assumed to be lsc, there is some positive real $\eta$ such that $f(x)\geq f(\bar{x})+\eta$ for all $x$ with $\|x-\bar{x}\|=\gamma$. Thus, whenever $\|x^{*}\|<\bar{\delta}:=\min\{\eta,\rho-f(\bar{x})\}/\gamma$ we have
$M_{\gamma}(x^{*}):=\mathop{\rm arg\,min}\{f(x)-\langle x^{*},x\rangle\,\mid\,x\in\bar{\cal B}_{\gamma}(\bar{x})\}=\mathop{\rm arg\,min}\{f(x)-\langle x^{*},x-\bar{x}\rangle\,\mid\,x\in\bar{\cal B}_{\gamma}(\bar{x})\}\subset{\cal B}_{\gamma}(\bar{x})$
implying $0\in\partial f(x)-x^{*}$, $x\in M_{\gamma}(x^{*})$, and
$f(x)-(\rho-f(\bar{x}))<f(x)-\langle x^{*},x-\bar{x}\rangle\leq f(\bar{x}),\ x\in M_{\gamma}(x^{*}).$
Therefore, for every $x^{*}\in{\cal B}_{\bar{\delta}}(0)$ and every $x\in M_{\gamma}(x^{*})$ we have $x^{*}\in\partial_{\rho}f(x)$ and lower semicontinuity of $f$ ensures that $M_{\gamma}(x^{*})$ is nonempty. By utilizing [22, Theorem 1.17], we can find some open neighborhood $\tilde{V}\subset{\cal B}_{\bar{\delta}}(0)$ of 00 such that $M_{\gamma}(x^{*})\subset U$, $x^{*}\in\tilde{V}$. It follows that for every $x^{*}\in V\cap\tilde{V}$ and every $x\in M_{\gamma}(x^{*})$ we have $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)$ implying $x=\vartheta(x^{*})$. Hence $M_{\gamma}$ coincides with $\vartheta$ on $V\cap\tilde{V}$ and is therefore single-valued and Lipschitz continuous with constant $\sigma^{-1}$ on $V\cap\tilde{V}$ verifying that $\bar{x}$ is a tilt-stable local minimizer for $f$ with modulus $\sigma^{-1}$. There remains to show that $f$ is prox-regular at $\bar{x}$ for 00. Choose $0<\epsilon<\min\{\gamma,\rho-f(\bar{x})\}$ such that ${\cal B}_{\epsilon}(\bar{x})\subset U$ and ${\cal B}_{\epsilon}(0)\subset\tilde{V}\cap V$. Then for every pair $(x,x^{*})\in\mathrm{gph}\,\partial_{f(\bar{x})+\epsilon}f\cap({\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}(0))$ we have $x=\vartheta(x^{*})=M_{\gamma}(x^{*})$ implying that for every $x^{\prime}\in{\cal B}_{\epsilon}(\bar{x})\subset{\cal B}_{\gamma}(\bar{x})$ there holds $f(x^{\prime})-\langle x^{*},x^{\prime}\rangle\geq f(x)-\langle x^{*},x\rangle$. From this inequality we easily deduce that $f$ is prox-regular at $\bar{x}$ for 00 with parameters $r=0$ and the chosen $\epsilon$. ∎
Remark 5.3.
Though Theorems 2.7, 5.1 and 5.2 provide a bunch of characterizations for tilt-stability and variational strong convexity for prox-regular functions, the list of equivalences can still be extended. E.g., by utilizing the relation (i)$\Leftrightarrow$(ii) from Theorem 2.7, one can easily add to Theorem 5.2 the following statement generalizing Theorem 3.2 of Mordukhovich and Nghia [13] (see also [1, Theorem 3.3 (3.)] in the paper of Drusvyatskiy and Lewis):
  1. (iv)
    There are neighborhoods $U$ of $\bar{x}$ and $V$ of ${\bar{x}^{\ast}}$ together with a real $\rho>f(\bar{x})$ such that the mapping $(\partial_{\rho}f)^{-1}$ admits a single-valued localization $\vartheta:V\to U$ around $({\bar{x}^{\ast}},\bar{x})$ and that for any pair $(x^{*},x)\in\mathrm{gph}\,\vartheta=\mathrm{gph}\,(\partial_{\rho}f)^{-1}\cap(V\times U)$ we have the uniform second-order growth condition
    $f(x^{\prime})\geq f(x)+\langle x^{*},x^{\prime}-x\rangle+\frac{\sigma}{2}\|x^{\prime}-x\|^{2}$
    whenever $x^{\prime}\in U$.
I am very grateful to one of the reviewers which asked for a comparison of Theorem 5.1 with [13, Theorem 3.2] of Mordukhovich and Nghia.. It motivated me to add statement (iii) to the equivalences in Theorem 5.2.
@@
    For every $s^{\prime}\in(-\infty,s)$, $f$ is variationally $s^{\prime}$-convex at $\bar{x}$ for ${\bar{x}^{\ast}}$.
@@
Choose $t>0$ such that $\sigma:=s+t>0$ and apply Theorem 5.1 to the function $\tilde{f}:=f+\frac{t}{2}\|\cdot-\bar{x}\|^{2}$. Since we can take $t$ arbitrarily large, the assertions follow from Proposition 3.2 and Lemma 4.2. ∎